Collaboration Tools

• The idea of the alliance is colloboration between various places.

• No collobration tools exisits yet.

• To date, the library repository is a failure.

• Reason why RNPL is not the direction to go to aid this colloboration

1. Routines such as clustering, interpolators, special types of differencing, etc. must be written independently of RNPL

2. RNPL does not supply an easy way to hook up the modules

3. RNPL is not WEB based

4. Must codes in the alliance would be written already in DAGH and

thus, would not change the interworkings of the code.

• Tool Kit approach 1. Look at how the “AVS” type interface helped colloboration.

2. We propose to build a similar interface.

3. By working on this in Java, we would create a user interface that will be available via the WWW.

4. Within this framework, we can use other collaborative tools to aid research.

5. Here interfaces will be designed for modules, and thus, modules can be swapped in and out to test the routines.

6. Clusters, etc. can be instantly swapped to see how the performance

is affected.

• Comonents of the toolkit 1. Module Generator

2. AMR driver

3. Interpolators

4. Outer boundary matching condition routines

5. Apparent Horizon trackers

6. Multigrid solvers

7. Stability modules

8. Graphical interfaces to visualization routines

9. VRML output